OpenStudy (anonymous):

Okay here's another but at least I got the GCF out 24r^3 + 166r^2 - 168r I factored out 4r to get: 4r (6r^2 + 29r - 42) but of course this is where I get stuck

OpenStudy (anonymous):

166/4 = 83/2

OpenStudy (anonymous):

I'm sorry wrote it wrong, ahhhhh, it's 116r not 166r

OpenStudy (anonymous):

original problem is 24r^3 + 116r^2 - 168r

OpenStudy (anonymous):

now u factorise the quadratic equation in the bracket...

OpenStudy (anonymous):

does the problem allow you to use the quadratic method of finding r?

OpenStudy (anonymous):

then i factored out 4r to get 4r(6r^2 + 29r - 42)

OpenStudy (anonymous):

ur factors are correct now factior the quadratic

myininaya (myininaya):

now looking at 6r^2+29r-42 we need to find two factors of 6(-42) that multiply to be 6(-42) and add up to be 29 6(-42)=36(-7) 36-7=29 so we have 6r^2+36r-7r-42 now factor by grouping

OpenStudy (anonymous):

hint: factors of -42 are +6 and -7

OpenStudy (anonymous):

i got (6r-7)(r+6)

OpenStudy (anonymous):

right

myininaya (myininaya):

so don't forget the 4r and you are done

OpenStudy (anonymous):

by jove you've got it!

OpenStudy (anonymous):

so it's 4r(6r-7)(r+6)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks guys so much